OpenStudy (anonymous):

The distance from R(3,-3,1) to the plane with the equation X-2y+6z=0 is 3. Determine all possible value(s) of A for which this is true.

OpenStudy (amistre64):

i can never recall the formulas for this so i tend to have to build something workable

OpenStudy (anonymous):

oh ok.. please don't use equation because it's not working for me:(

OpenStudy (amistre64):

pull the normal vector from the plane and define the line using it and the given point in parametric form use the parametric forms to define the point on the plane where it intersects (solving for t), then plugging in t into the parametric equations defines the second point ... the distance between them is therefore apparent

OpenStudy (amistre64):

all possible values of "A"?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

The distance from R(3,-3,1) to the plane with the equation X-2y+6z=0 is 3 there is no A in the post ....

OpenStudy (anonymous):

oh!!!! my bad.. Ax-2y+6z..

OpenStudy (psymon):

It's assumed as a part of the equation for a plane usually :P

OpenStudy (amistre64):

normal vector is therefore: (a,-2,6), anchored to (3,-3,1) im thinking at first glance that this is a sphere centered at the given point :)

OpenStudy (anonymous):

give me a a sec. i'll try

OpenStudy (psymon):

*watches and learns*

OpenStudy (amistre64):

x = 3 + at y = -3 - 2t z = 1 + 6t ax-2y+6z = 0 a(3+at)-2(-3-2t)+6(1+6t) = 0 3a+a^2 t + 6 + 4t + 6 +36t = 0 (40+a^2)t + 12+3a = 0 t = -(12+3a)/(40+a^2)

OpenStudy (amistre64):

the point on the plane is therefore: x = 3 - a(12+3a)/(40+a^2) y = -3 + 2(12+3a)/(40+a^2) z = 1 - 6(12+3a)/(40+a^2) and using the distance formula between (3,-3,1) and this point in "a"; equaled to 3 should solve for all a

OpenStudy (anonymous):

WOW.. thanks a lot.. (^_^)

OpenStudy (anonymous):

@Psymon you understand??

OpenStudy (amistre64):

formulas would most likely make the computations plug and chug .. but the process should still amount to the same resutls :)

OpenStudy (psymon):

Yeah, looks like a similar procedure to what we were doing. Im at work, sorry, haha. I work as a math tutor, Im just not as high of a level as Id like xD

OpenStudy (amistre64):

moving the setup to the origin should subtract out the point parts leaving us with the "a" values ... x = - a(12+3a)/(40+a^2) y = 2(12+3a)/(40+a^2) z = - 6(12+3a)/(40+a^2) 9 = x^2 + y^2 + z^2

OpenStudy (psymon):

And if youre like me, I actually save what is said on these problems xD

OpenStudy (amistre64):

i never save anything, which means i have to reconstruct it on the fly ;)

OpenStudy (psymon):

Well if I remember I had a problem doneand worked out, I can refer to the steps and what the person who helped me did.